# Thermodynamically consistent treatment of repulsive corrections in HRG

### v1

@@ I Introduction
(2) the thermodynamic properties are highly sensitive to the size of the hadrons. Unfortunately, excluded volume of each and every hadron species is not known yet. Although there is a consensus regarding the value of the pion radius to be of the order of 0.3 $fm$, no conclusive results exist for most of the other hadrons.
@@ II.2 Conceptual framework to facilitate the calculation
Here, $g_{i}$ is the spin degeneracy of hadronic species $i$; $m_{i}$ is the mass, $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ are the modified Bessel functions. The Pressure $(P)$ of system is evaluated in the original quantum formulation.
@@ III Results
To show the importance of the proper treatment of the non-zero chemical potential term in the effective free particle picture, we compare the conserved charge susceptibilities under two different scenario: (1) Model I with the chemical potential term, arising due to repulsive interactions, absorbed in the individual particle energy level, as described in the previous section of this paper. (2) Model II with the traditional treatment of the effective chemical potential, where no equivalent effective particle picture has been considered. We show the results for pion radius $r_{\pi}$ = 0.2 $fm$ and $\mathcal{A}$=0.5 and investigate the conserved charge susceptibilities at zero chemical potentials. In this work, we use the QMHRG2020 list of hadrons Bollweg et al. (2021b), which includes all the confirmed (3-and 4-star baryon resonances) and the unconfirmed (1- and 2-star baryon resonances as well as mesons not listed in the PDG 2020 summary tables) hadrons. Additionally, this list also includes quark-model states in the strange and non-strange baryon sectors. All the baryon number, electric charge and strangeness susceptibilities, except the fourth order electric charge susceptibility, are well reproduced within the ”Model I” except $\chi_{Q}^{4}$. It is worth mentioning that the lattice calculations of the fourth-order electric charge susceptibility exhibits comparatively larger lattice uncertainties. Agreement in the higher order strange sector is also not good which may be due to the undiscovered high mass strange hadrons. The charge susceptibilities of crossed order containing baryon number as one of the conserved charge also show good agreement with the lattice data.
@@ IV Summary
In this work, we reformulate the treatment of density-dependent chemical potential shifts appearing in excluded-volume implementations by mapping the quantum statistics picture to an equivalent classical statistics picture and assigning a common ground state energy level for all the hadron species. Thermodynamic consistency in the classical statistics picture is maintained by equating the number densities in both pictures. Assuming the hadrons as classical liquid drops, a phenomenological formula for the hadron radius varying as power law over mass has been derived. We find that the resulting model satisfactorily reproduces the lattice data for all the lower order conserved charge susceptibilities simultaneously with only two model parameters: pion radius $R_{\pi}=0.2\ fm$ and $\mathcal{A}=0.5$.

### v2

@@ I Introduction
(2) the thermodynamic properties are highly sensitive to the size of the hadrons. Unfortunately, excluded volume of each and every hadron species is not known yet. Although there is a consensus regarding the value of the pion radius to be of the order of 0.2 $fm$, no conclusive results exist for most of the other hadrons.
@@ II.2 Conceptual framework to facilitate the calculation
Here, $g_{i}$ is the spin degeneracy of hadronic species $i$; $m_{i}$ is the mass, $\mathcal{K}_{2}$ is the modified Bessel functions. The Pressure $(P)$ of system is evaluated in the original quantum formulation.
@@ III Results
To show the importance of the proper treatment of the non-zero chemical potential term in the effective free particle picture, we compare the conserved charge susceptibilities under two different scenario: (1) Model I with the chemical potential term, arising due to repulsive interactions, absorbed in the individual particle energy level, as described in the previous section of this paper. (2) Model II with the traditional treatment of the effective chemical potential, where no equivalent effective particle picture has been considered. We show the results for pion radius $r_{\pi}$ = 0.2 $fm$ and $\mathcal{A}$=3 and investigate the conserved charge susceptibilities at zero chemical potentials. In this work, we use the QMHRG2020 list of hadrons Bollweg et al. (2021b), which includes all the confirmed (3-and 4-star baryon resonances) and the unconfirmed (1- and 2-star baryon resonances as well as mesons not listed in the PDG 2020 summary tables) hadrons. Additionally, this list also includes quark-model states in the strange and non-strange baryon sectors. All the baryon number, electric charge and strangeness susceptibilities, except the fourth order electric charge susceptibility, are well reproduced within the ”Model I” except $\chi_{Q}^{4}$. It is worth mentioning that the lattice calculations of the fourth-order electric charge susceptibility exhibits comparatively larger lattice uncertainties. Agreement in the higher order strange sector is also not good which may be due to the undiscovered high mass strange hadrons. The charge susceptibilities of crossed order containing baryon number as one of the conserved charge also show good agreement with the lattice data.
@@ IV Summary
In this work, we reformulate the treatment of density-dependent chemical potential shifts appearing in excluded-volume implementations by mapping the quantum statistics picture to an equivalent classical statistics picture and assigning a common ground state energy level for all the hadron species. Thermodynamic consistency in the classical statistics picture is maintained by equating the number densities in both pictures. Assuming the hadrons as classical liquid drops, a phenomenological formula for the hadron radius varying as power law over mass has been derived. We find that the resulting model satisfactorily reproduces the lattice data for all the lower order conserved charge susceptibilities simultaneously with only two model parameters: pion radius $R_{\pi}=0.2\ fm$ and $\mathcal{A}=3$.